OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

y = mx + b m = slope y = -3x + 2 Remember, parallel lines have THE SAME slope. Okay, you can already eliminate two choices, which ones are those?

OpenStudy (anonymous):

a and c?

OpenStudy (igreen):

Yes! Because their slopes are 1/3, and you need slopes with -3.

OpenStudy (igreen):

So B and D are our choices. We can plug in (-1, 6) into both of them to see which one it is. y = -3x + 3 Plug in (-1, 6): 6 = -3(-1) + 3 Can you finish that?

OpenStudy (anonymous):

I'm still a little confused

OpenStudy (igreen):

Okay, well we have to multiply -3 and -1, can you do that?

OpenStudy (anonymous):

3?

OpenStudy (igreen):

Yep. That gives us 6 = 3 + 3. Now add that and what do you get?

OpenStudy (anonymous):

6

OpenStudy (igreen):

Yep, so it gives you 6 = 6, and that's true. So y = -3x + 3 is the correct answer.

OpenStudy (anonymous):

I have 1 last qestion thats it will you help?

OpenStudy (igreen):

Sure.